### v2

@@ Abstract
Large language model (LLM) queries are predominantly processed by frontier models in centralized cloud infrastructure. Rapidly growing demand strains this paradigm, and cloud providers struggle to scale infrastructure at pace. Two advances create an opportunity to rethink this paradigm: small, local LMs ($\leq$ 20B active parameters) now achieve competitive performance to frontier models on many tasks, and local accelerators (e.g., Apple M4 Max) can host these models at interactive latencies. This raises the question: can local inference viably redistribute demand from centralized infrastructure? Answering this requires measuring both whether local LMs can accurately answer real-world queries and whether they can do so efficiently enough to be practical on power-constrained devices (i.e., laptops). We propose intelligence per watt (IPW), task accuracy divided by unit of power, as a unified metric for assessing both the capability and efficiency of local inference across model-accelerator configurations. We conduct a large-scale empirical study across 20+ state-of-the-art local LMs, 8 hardware accelerators (local and cloud), and a representative subset of LLM traffic: 1M real-world single-turn chat and reasoning queries. For each query, we measure accuracy (local LM win rate against frontier models), energy consumption, latency, and power. Our analysis reveals three key findings. First, that local LMs can successfully answer $88.7\%$ of single-turn chat and reasoning queries with accuracy varying by domain. Second, longitudinal analysis from 2023-2025 shows progress in local inference viability: IPW improved $5.3\times$, driven by both algorithmic advances and accelerator improvements, with locally-serviceable query coverage increasing from 23.2% to 71.3%. Third, local accelerators achieve at least 1.4× lower IPW than cloud accelerators running identical models, revealing significant headroom for local accelerator optimization These findings demonstrate that local inference can meaningfully redistribute demand from centralized infrastructure for a substantial subset of queries, with IPW serving as the critical metric for tracking this transition. We release our IPW profiling harness to enable intelligence-per-watt benchmarking as local LMs and accelerators evolve.
@@ 5.2 How Intelligence Efficient is Local Inference?
SOTA
Local Model
Mixtral-8x7B-v0.1 Llama-3.1-8B-Instruct GPT-OSS-120B
SOTA
Accelerator
NVIDIA Quadro
RTX 6000
NVIDIA RTX
6000 Ada
Apple
M4 Max
Success Rate
$23.2\pm 1.9\%$ $48.7\pm 2.7\%$ $71.3\pm 2.2\%$
Intelligence
per Watt
$(7.92\pm 0.32)$
$\times 10^{-4}$
$(1.80\pm 0.21)$
$\times 10^{-3}$
$(4.18\pm 0.53)$
$\times 10^{-3}$
YoY Efficiency
Gain
2.27× 2.32×
@@
Intelligence
per Watt
$(1.40\pm 0.38)$
$\times 10^{-3}$
$(1.63\pm 0.20)$
$\times 10^{-3}$
$(1.69\pm 0.31)$
$\times 10^{-3}$
$(1.97\pm 0.24)$
$\times 10^{-3}$
@@
Intelligence
per Watt
$(1.95\pm 0.14)$
$\times 10^{-3}$
$(2.27\pm 0.18)$
$\times 10^{-3}$
$(2.35\pm 0.09)$
$\times 10^{-3}$
$(2.75\pm 0.14)$
$\times 10^{-3}$
@@
Intelligence
per Watt
$(3.51\pm 0.43)$
$\times 10^{-3}$
@@
Intelligence
per Joule
$(3.80\pm 0.40)$
$\times 10^{-5}$
$(3.51\pm 0.38)$
$\times 10^{-5}$
$(4.38\pm 0.31)$
$\times 10^{-5}$
$(4.23\pm 0.45)$
$\times 10^{-5}$
@@
Intelligence
per Joule
$(8.71\pm 0.60)$
$\times 10^{-5}$
$(5.91\pm 0.42)$
$\times 10^{-5}$
$(7.34\pm 0.51)$
$\times 10^{-5}$
$(6.78\pm 0.47)$
$\times 10^{-5}$
@@
Intelligence
per Joule
$(2.27\pm 0.30)$
$\times 10^{-4}$
$(3.12\pm 0.38)$
$\times 10^{-4}$
@@ 8 Acknowledgements
We thank Simran Arora, Bradley Brown, Mayee Chen, Owen Dugan, Sabri Eyuboglu, Neel Guha, Simon Guo, Jordan Juravsky, Jerry Liu, Benjamin Spector, Shayan Talaei, and Dan Zhang for their constructive feedback during the composition of the paper. We would also like to thank our collaborators at the Stanford Artificial Intelligence Laboratory (SAIL) and TogetherAI.
@@ 36 Category:
Domain
WC
Count
WC
WC
Solv %
NR
Count
NR
NR
Solv %
@@
Arts, design, sports, entertainment,
and media
235,658 47.1 98.7 2,648 0.5 52.9
Life, physical, and
social science
28,079 5.6 98.8 180,065 36.0 60.5
@@
Healthcare practitioners
and technical
8,905 1.8 98.1 1,851 0.4 66.3
Office and administrative
support
6,959 1.4 91.6 25 0.0 48.0
@@
Transportation and
material moving
1,890 0.4 95.1 3,914 0.8 53.0
@@
Food preparation and
serving related
3,308 0.7 98.3 507 0.1 61.9
@@
Installation, maintenance,
and repair
954 0.2 97.1 2,037 0.4 56.2
@@
Building grounds cleaning
and maintenance
278 0.1 100.0 70 0.0 72.9
@@ 94. Technical precision
Size Threshold ($\leq$)
Qwen
+ GPT-OSS
Qwen
Qwen
+ GPT-OSS
Qwen
Qwen
+ GPT-OSS
Qwen
@@
Size Threshold ($\leq$)
Qwen
+ GPT-OSS
Qwen
Qwen
+ GPT-OSS
Qwen
Qwen
+ GPT-OSS
Qwen

### v3

@@ Abstract
Large language model (LLM) queries are predominantly processed by frontier models in centralized cloud infrastructure. Rapidly growing demand strains this paradigm, and cloud providers struggle to scale infrastructure at pace. Two advances create an opportunity to rethink this paradigm: small, local LMs ($\leq$ 20B active parameters) now achieve competitive performance to frontier models on many tasks, and local accelerators (e.g., Apple M4 Max) can host these models at interactive latencies. This raises the question: can local inference viably redistribute demand from centralized infrastructure? Answering this requires measuring both whether local LMs can accurately answer real-world queries and whether they can do so efficiently enough to be practical on power-constrained devices (i.e., laptops). We propose intelligence per watt (IPW), task accuracy divided by unit of power, as a unified metric for assessing both the capability and efficiency of local inference across model-accelerator configurations. We conduct a large-scale empirical study across 20+ state-of-the-art local LMs, 8 hardware accelerators (local and cloud), and a representative subset of LLM traffic: 1M real-world single-turn chat and reasoning queries. For each query, we measure accuracy (local LM win rate against frontier models), energy consumption, latency, and power. Our analysis reveals three key findings. First, local LMs can successfully answer $88.7\%$ of single-turn chat and reasoning queries with accuracy varying by domain. Second, longitudinal analysis from 2023-2025 shows progress in local inference viability: IPW improved $5.3\times$, driven by both algorithmic advances and accelerator improvements, with locally-serviceable query coverage increasing from 23.2% to 71.3%. Third, local accelerators achieve at least 1.4× lower IPW than cloud accelerators running identical models, revealing significant headroom for local accelerator optimization These findings demonstrate that local inference can meaningfully redistribute demand from centralized infrastructure for a substantial subset of queries, with IPW serving as the critical metric for tracking this transition. We release our IPW profiling harness to enable intelligence-per-watt benchmarking as local LMs and accelerators evolve: https://github.com/HazyResearch/intelligence-per-watt.
@@ 5.2 How Intelligence Efficient is Local Inference?
SOTA Local Model Mixtral-8x7B-v0.1 Llama-3.1-8B-Instruct GPT-OSS-120B
SOTA Accelerator NVIDIA Quadro RTX 6000 NVIDIA RTX 6000 Ada Apple M4 Max
Success Rate $23.2\pm 1.9\%$ $48.7\pm 2.7\%$ $71.3\pm 2.2\%$
Intelligence per Watt $(7.92\pm 0.32)$ $\times 10^{-4}$ $(1.80\pm 0.21)$ $\times 10^{-3}$ $(4.18\pm 0.53)$ $\times 10^{-3}$
YoY Efficiency Gain 2.27× 2.32×
@@
Intelligence per Watt $(1.40\pm 0.38)$ $\times 10^{-3}$ $(1.63\pm 0.20)$ $\times 10^{-3}$ $(1.69\pm 0.31)$ $\times 10^{-3}$ $(1.97\pm 0.24)$ $\times 10^{-3}$
@@
Intelligence per Watt $(1.95\pm 0.14)$ $\times 10^{-3}$ $(2.27\pm 0.18)$ $\times 10^{-3}$ $(2.35\pm 0.09)$ $\times 10^{-3}$ $(2.75\pm 0.14)$ $\times 10^{-3}$
@@
Intelligence per Watt $(3.51\pm 0.43)$ $\times 10^{-3}$
@@
Intelligence per Joule $(3.80\pm 0.40)$ $\times 10^{-5}$ $(3.51\pm 0.38)$ $\times 10^{-5}$ $(4.38\pm 0.31)$ $\times 10^{-5}$ $(4.23\pm 0.45)$ $\times 10^{-5}$
@@
Intelligence per Joule $(8.71\pm 0.60)$ $\times 10^{-5}$ $(5.91\pm 0.42)$ $\times 10^{-5}$ $(7.34\pm 0.51)$ $\times 10^{-5}$ $(6.78\pm 0.47)$ $\times 10^{-5}$
@@
Intelligence per Joule $(2.27\pm 0.30)$ $\times 10^{-4}$ $(3.12\pm 0.38)$ $\times 10^{-4}$
@@ 8 Acknowledgements
We thank Simran Arora, Kush Bhatia, Bradley Brown, Rishi Bommasani, Mayee Chen, Owen Dugan, Sabri Eyuboglu, Neel Guha, Simon Guo, Braden Hancock, Jordan Juravsky, Omar Khattab, Andy Konwinski, Jerry Liu, Kunle Olukotun, Christopher Rytting, Benjamin Spector, Shayan Talaei, and Dan Zhang for their constructive feedback during the composition of the paper. We would also like to thank our collaborators at the Stanford Artificial Intelligence Laboratory (SAIL) and TogetherAI.
@@ 36 Category:
Domain WC Count WC % WC Solv % NR Count NR % NR Solv %
@@
Arts, design, sports, entertainment, and media 235,658 47.1 98.7 2,648 0.5 52.9
Life, physical, and social science 28,079 5.6 98.8 180,065 36.0 60.5
@@
Healthcare practitioners and technical 8,905 1.8 98.1 1,851 0.4 66.3
Office and administrative support 6,959 1.4 91.6 25 0.0 48.0
@@
Transportation and material moving 1,890 0.4 95.1 3,914 0.8 53.0
@@
Food preparation and serving related 3,308 0.7 98.3 507 0.1 61.9
@@
Installation, maintenance, and repair 954 0.2 97.1 2,037 0.4 56.2
@@
Building grounds cleaning and maintenance 278 0.1 100.0 70 0.0 72.9
@@ 94. Technical precision
Size Threshold ($\leq$) Qwen + GPT-OSS Qwen Qwen + GPT-OSS Qwen Qwen + GPT-OSS Qwen
@@
Size Threshold ($\leq$) Qwen + GPT-OSS Qwen Qwen + GPT-OSS Qwen Qwen + GPT-OSS Qwen